### v1

@@ 1 Introduction
With the rise of Large Language Models (LLMs), generative pre-training has become increasingly popular. Representations learned via next token prediction improves the performance when transferred to a diverse set of downstream tasks (Radford et al., 2018; 2021; Brown et al., 2020; Raffel et al., 2023). Beyond learning representations, the model directly generates language, acting as a user interface and allowing for interactive adjustments (Liu et al., 2021). The likelihood-based pre-training objective enables us to investigate scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022). These scaling laws predict how a model’s pre-training loss relates to its capacity and the amount of training data. Generally, we anticipate that achieving a lower pre-training loss leads to superior performance on downstream tasks.
In vision, however, representation learning and image generation often use separate techniques. For learning representations, methods such as contrastive learning van den Oord et al. (2019); Chen et al. (2020b); He et al. (2019), distillation-based self-supervised learning Grill et al. (2020); Caron et al. (2021) and masked image modelling (MIM) He et al. (2021); Bao et al. (2022) are widely used. Despite their strengths in learning robust visual and cross-modal Radford et al. (2021) representations, as well as their efficient use of model capacity, these methods lack generation capabilities. Furthermore, the pre-training loss, influenced by the difficulty of the pre-training task, does not serve as a reliable indicator of performance on downstream tasks.
@@
  • MSE and diffusion objectives: We observe that training on MSE loss alone yields strong performance. Incorporating a denoising patch decoder further enhances representation quality and generative ability, especially in larger models with extended training and optimized noise schedules.
@@ 2 Related Works
Generative Pre-Training. In the vision domain, earlier attempts of using generative models for representation learning includes Variational Autoencoders (VAE) Kingma & Welling (2022); Rezende et al. (2014); Higgins et al. (2017); van den Oord et al. (2018) and GANs (Donahue & Simonyan, 2019). With the success of GPT (Radford et al., 2018; 2021; Brown et al., 2020) in NLP, generative pre-training attracts renewed attention. Image-GPT (Chen et al., 2020a) adapts the GPT model for pre-training on images.
@@ 3.2 Training Objective
$\mathcal{L}_{MSE}(\theta;\mathcal{D})\propto\sum_{\mathcal{D}}\sum_{t=1}^{T}\lVert\hat{x}_{t}-x_{t}\rVert^{2}$
This can be interpreted as modelling $x_{t}$ with a Gaussian distribution centered at $\hat{x}_{t}$ with a constant variance. Despite the probabilistic interpretation of MSE loss, it is rarely used in state-of-the-art generative models, because the unimodal assumption of the Gaussian distribution leads to blurry images.
@@
3.3 Rotary Positional Embedding for Images
Relative positional encoding, notably rotary positional embedding (RoPE) Su et al. (2023), is commonly used in NLP. RoPE is particularly popular due to its ease of implementation and overall improvement to performance. However, in vision, the common practice is to use absolute positional encoding or learnable positional embedding due to the low accuracy/compute trade-off or the speed of training of the relative positional encoding Dosovitskiy et al. (2021). We propose to extend RoPE to higher dimensional positions by decomposing and applying RoPE separately for each dimension.
To be more precise, let $m=(m_{x},m_{y})$ and $n=(n_{x},n_{y})$ be the $x$ and $y$ coordinates of two patches. In the self-attention layer, let $z^{k}_{m}=W_{k}x_{m}$ be the activation after linear key projection for patch at location $m$; $z^{q}_{n}=W_{q}x_{n}$ be the activation linear query projection at location $n$. The rotation matrix $R$ is a block diagonal matrix where each block uses a different frequency and applies to either $x$ or $y$ coordinate. For the simplified case where the channel dimension of $z^{k}$ and $z^{q}$ is 4, we construct a rotation matrix with one frequency for each coordinate:
$k_{m}=R_{\theta}(m)z^{k}_{m},\quad q_{n}=R_{\theta}(n)z^{q}_{n}$
$k_{m},q_{n}\right>=(z^{k}_{m})^{T}R_{\theta}(m)^{T}R_{\theta}(n)z^{q}_{n}$
The decomposed RoPE is easy to implement, requires minimal changes relative to 1D RoPE, and readily extends to higher-dimensional coordinates. However, splitting features by coordinate dimensions reduces the frequency coverage per dimension. We did not encounter this issue with 2D data, but if it arises, consider projecting activations into a higher-dimensional space before applying the rotation matrix.
3.4 Diffusion
@@
$\mathcal{L}_{\text{SIMPLE}}$ $(x_{t};x_{<t},f,g)=\mathbb{E}_{s,\epsilon}\left[\lVert x^{0}_{t}-g(x^{s}_{t},f(x_{<t}))\rVert^{2}\right]$
@@ 4 Experiments
The input image has a resolution of 224 $\times$ 224. Spatial augmentation (random cropping, resizing and horizontal flipping) is applied during pre-training. The initialization scheme and optimization follow the MAE recipe He et al. (2021). Pre-training uses AdamW (Loshchilov & Hutter, 2019) for optimization with learning rate using cosine decay and 40 epochs warm-up. The full list of hyper-parameters can be found in Table 7 in the Appendix.
@@
Due to the lack of bottleneck, the network learns a representation that is more distributed across the network. As a result, fine-tuning is a better evaluation protocol. However, in Appendix B, we provide a more in-depth study of linear evaluation results. We find that, although no explicit bottleneck is imposed, the best performing layer is situated roughly in the middle of the Transformer stacks.
@@ 4.1 Main Properties
We compare the decomposed RoPE (2D RoPE) with NoPE Kazemnejad et al. (2023), absolute positional encodings and learnable positional embedding. NoPE doesn’t apply any positional encodings. Absolute positional encodings uses the sine and cosine functions of various frequencies to encode the 2D position (Vaswani et al., 2023; He et al., 2021). Learnable positional embedding is randomly initialized and adapted with the training. In addition, we compare 2D RoPE which uses the inductive bias of spatial coordinates with the vanilla RoPE (1D RoPE). To evaluate or fine-tune models on a higher resolution, we interpolate the positions when using RoPE.
@@
We take a closer look at the patch decoders and make architectural comparison between a linear patch decoder, a residual MLP and a Transformer patch decoder with varying depth. The Transformer patch decoder only applies for the denoising patch decoder: the sequence consists of two tokens - the output of the backbone and the embedding of the noise corrupted patch, both at the same autoregressive step $t$. No causal attention masking is applied for the Transformer patch decoder. In addition to the architectural choices, we experiment with whether or not to condition on the diffusion step $\gamma$. When conditioned on $\gamma$, it’s added as a channel to the noise corrupted image patch and linearly projected to the embedding space.
@@ 4.2 Comparison to Previous Results
Table 4: ImageNet top-1 Accuracy Comparison
Backbone
Category Method ViT-B ViT-L ViT-H Others
Contrastive DINO 1 82.8 - - -
MoCo v3 2 83.2 84.1 - -
Masked Pred. BeiT 3 83.2 85.2 - -
MAE 4 83.6 85.9 86.9 -
Generative Image-GPT 5 - - - 72.6
DARL (MSE) 82.7 84.7 85.5 -
DARL (DIFF) 81.9 84.9 85.9 -
  • 1
    Caron et al. (2021)
  • 2
    He et al. (2019)
  • 3
    Bao et al. (2022)
  • Tokenizer is trained on a much larger custom dataset Ramesh et al. (2021).
  • 4
    He et al. (2021)
  • 5
    Chen et al. (2020a)
@@ A.2 Supervised Baseline
The supervised baseline is similar to the ViT Dosovitskiy et al. (2021) implementation with the following differences:
@@
A.3 Two-Stream Architecture
@@ A.4 VTAB
Appendix B Linear Evaluation
@@
(a) Training Epochs
(b) Network Depth
Figure 8: ImageNet top-1 accuracy with linear evaluation protocol. 7(a)) Linear performance increases with longer training schedule. Models are ViT-L16 pre-trained with diffusion objective. 7(b)) Linear performance varies with the layer depth. Layer depth is normalized to [0, 1]. Models are pre-trained with diffusion objective.
@@ Appendix C Visualizations for Patch Ordering
Figure 9 compares the patch reconstruction error at each postion in the sequence between different fixed order strategies. Figure 10 and Figure 11 show the visualization and the average per-patch reconstruction errors for nested raster order. Figure 12 and Figure 13 show the visualization and the average per-patch reconstruction errors for round-robin order. See Table 6 for details.
@@ Appendix E Diffusion
We use the mathematical framework established by DDPM Ho et al. (2020). Many of the analysis can be found in DDPM paper. We present here for the completeness of our analysis. To simplify notation, we use $x$ as the image or individual patch and subscript $t$ for the diffusion process.

### v2

@@ 1 Introduction
With the rise of Large Language Models (LLMs), generative pre-training has become increasingly popular. Representations learned via next token prediction improves the performance when transferred to a diverse set of downstream tasks (Radford et al., 2018, 2021; Brown et al., 2020; Raffel et al., 2023). Beyond learning representations, the model directly generates language, acting as a user interface and allowing for interactive adjustments (Liu et al., 2021). The likelihood-based pre-training objective enables us to investigate scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022). These scaling laws predict how a model’s pre-training loss relates to its capacity and the amount of training data. Generally, we anticipate that achieving a lower pre-training loss leads to superior performance on downstream tasks.
In vision, however, representation learning and image generation often use separate techniques. For learning representations, methods such as contrastive learning (van den Oord et al., 2019; Chen et al., 2020b; He et al., 2019), distillation-based self-supervised learning (Grill et al., 2020; Caron et al., 2021) and masked image modelling (MIM) (He et al., 2021; Bao et al., 2022) are widely used. Despite their strengths in learning robust visual and cross-modal (Radford et al., 2021) representations, as well as their efficient use of model capacity, these methods lack generation capabilities. Furthermore, the pre-training loss, influenced by the difficulty of the pre-training task, does not serve as a reliable indicator of performance on downstream tasks.
@@
  • MSE and diffusion objectives: We observe that training on MSE loss alone yields strong performance. Incorporating a denoising patch decoder further enhances representation quality and generative ability, especially in larger models with extended training and optimized noise schedules. Denoising is also beneficial when using large patch sizes.
@@ 2 Related Works
Generative Pre-Training. In the vision domain, earlier attempts of using generative models for representation learning includes Variational Autoencoders (VAE) (Kingma & Welling, 2022; Rezende et al., 2014; Higgins et al., 2017; van den Oord et al., 2018) and GANs (Donahue & Simonyan, 2019). With the success of GPT (Radford et al., 2018, 2021; Brown et al., 2020) in NLP, generative pre-training attracts renewed attention. Image-GPT (Chen et al., 2020a) adapts the GPT model for pre-training on images.
@@ 3.2 Training Objective
$\mathcal{L}_{MSE}(\theta;\mathcal{D})\propto\sum_{\mathcal{D}}\sum_{t=1}^{T}\lVert f(x_{<t})-x_{t}\rVert^{2}$
$f(x_{<t})$ is output of the Transformer. This can be interpreted as modelling $x_{t}$ with a Gaussian distribution centered at $f(x_{<t})$ with a constant variance. Despite the probabilistic interpretation of MSE loss, it is rarely used in state-of-the-art generative models, because the unimodal assumption of the Gaussian distribution leads to blurry images.
@@
3.3 Patch Diffusion
@@
$\mathcal{L}_{\text{SIMP}}$ $(x_{t};x_{<t},f,g)=\mathbb{E}_{\gamma\sim\mathcal{B},\epsilon}\left[\lVert x^{0}_{t}-g(x^{s(\gamma)}_{t},f(x_{<t}))\rVert^{2}\right]$
3.4 Rotary Positional Embedding for Images
While rotary positional embedding (RoPE) (Su et al., 2023) is widely used in NLP, its application in vision has been limited due to perceived accuracy/compute trade-offs and slower training compared to absolute or learnable positional encoding (Dosovitskiy et al., 2021). We extend RoPE to higher-dimensional positions by applying it separately to each dimension, addressing these limitations.
Let $m=(m_{x},m_{y})$ and $n=(n_{x},n_{y})$ represent the coordinates of two patches. In the self-attention layer, let $z^{k}_{m}=W_{k}x_{m}$ and $z^{q}_{n}=W_{q}x_{n}$ be the key and query projections for patches at locations $m$ and $n$, respectively. The rotation matrix $R$ is a block diagonal matrix with different frequencies for each coordinate ($x$ or $y$). For a simplified case with 4 channels in $z^{k}$ and $z^{q}$, the rotation matrix with one frequency per coordinate is:
$k_{m}=R_{\theta}(m)z^{k}_{m},\quad q_{n}=R_{\theta}(n)z^{q}_{n}$
$k_{m},q_{n}\right>=(z^{k}_{m})^{T}R_{\theta}(m)^{T}R_{\theta}(n)z^{q}_{n}$
The decomposed RoPE is easy to implement, requires minimal changes relative to 1D RoPE, and readily extends to higher-dimensional coordinates. However, splitting features by coordinate dimensions reduces the frequency coverage per dimension. We did not encounter this issue with 2D data, but if it arises, consider projecting activations into a higher-dimensional space before applying the rotation matrix. The extension of RoPE described here has been independently proposed and implemented by the pytorch library rotary-embedding-torch.
@@ 4 Experiments
The input image has a resolution of 224 $\times$ 224. Spatial augmentation (random cropping, resizing and horizontal flipping) is applied during pre-training. The initialization scheme and optimization follow the MAE recipe (He et al., 2021). Pre-training uses AdamW (Loshchilov & Hutter, 2019) for optimization with learning rate using cosine decay and 40 epochs warm-up. The full list of hyper-parameters can be found in Table 8 in the Appendix.
@@
Due to the lack of bottleneck, the network learns a representation that is more distributed across the network. As a result, fine-tuning is a better evaluation protocol. However, in Section B.1, we provide a more in-depth study of linear evaluation results. We find that, although no explicit bottleneck is imposed, the best performing layer is situated roughly in the middle of the Transformer stacks.
@@ 4.1 Main Properties
We compare the decomposed RoPE (2D RoPE) with NoPE (Kazemnejad et al., 2023), absolute positional encodings and learnable positional embedding. NoPE doesn’t apply any positional encodings. Absolute positional encodings uses the sine and cosine functions of various frequencies to encode the 2D position (Vaswani et al., 2023; He et al., 2021). Learnable positional embedding is randomly initialized and adapted with the training. In addition, we compare 2D RoPE which uses the inductive bias of spatial coordinates with the vanilla RoPE (1D RoPE). To evaluate or fine-tune models on a higher resolution, we interpolate the positions when using RoPE.
@@
We compare between a linear patch decoder, a residual MLP and a Transformer patch decoder with varying depth. The Transformer patch decoder only applies for the denoising patch decoder: the patch decoding sequence consists of two tokens - the output of the backbone and the embedding of the noise corrupted patch, both at the same autoregressive step $t$. No causal attention masking is applied for the Transformer patch decoder. In addition to the architectural choices, we experiment with whether or not to condition on the diffusion step $\gamma$. When conditioned on $\gamma$, it’s added as a channel to the noise corrupted image patch and linearly projected to the embedding space.
@@
Concurrent work (Chen et al., 2024) prefers a different noise schedule than ours, likely due to their focus on denoising as the pre-text task. Our framework combines autoregressive prediction with denoising, suggesting that autoregressive prediction is the primary driver of representation learning, with denoising providing additional benefits under specific conditions. Section B.5 provides more details.
@@ 4.2 Comparison to Previous Results
Table 5: ImageNet top-1 Accuracy Comparison
Backbone
Category Method ViT-B ViT-L ViT-H Others
Contrastive DINO 1 82.8 - - -
MoCo v3 2 83.2 84.1 - -
Masked Pred. BeiT 3 83.2 85.2 - -
MAE 4 83.6 85.9 86.9 -
Generative Image-GPT 5 - - - 72.6
DARL (MSE) 82.7 84.7 85.5 -
DARL (DIFF) 81.9 84.9 85.9 -
  • 1
    Caron et al. (2021)
  • 2
    He et al. (2019)
  • 3
    Bao et al. (2022)
  • Tokenizer is trained on a much larger custom dataset (Ramesh et al., 2021).
  • 4
    He et al. (2021)
  • 5
    Chen et al. (2020a)
@@
Table 6: COCO Object Detection and Segmentation
SUP MAE † DARL (DIFF)
AP 54.1 57.2 56.4
mAP 45.8 48.6 48.0
ViT-L16 backbone fine-tuned with ViTDet architecture and Cascade Mask R-CNN detectors. †MAE results are reproduced from our own codebase.
@@ 4.3 Transfer Experiments
Object detection and segmentation.
To assess the transfer capability of DARL in tasks other than classification, we fine-tune ViTDet (Li et al., 2022) with Cascade Mask R-CNN (Cai & Vasconcelos, 2017) as detector heads. Other implementations details are in Section A.3.
Table 6 confirms that, similar to ImageNet classification, DARL outperforms the supervised pre-training baseline and performs closely to the strong MAE baseline. ViTDet architecture uses a simple feature pyramid design which takes only the last layer activation of the Transformer. This design was developed with supervisely trained models and encoder-only representations (bottleneck at the last layer of the Transformer stack). It could be sub-optimal for generative pre-trained models such as ours. We leave this for future research endeavors.
@@ A.2 Supervised Baseline
The supervised baseline is similar to the ViT (Dosovitskiy et al., 2021) implementation with the following differences:
@@
A.3 COCO Object Detection and Segmentation
We simply initialize the weight of the ViT-L16 backbone with weight after supervised, MAE and DARL pre-training. Our implementation follows that of ViTDet in Li et al. (2022) and Cascade Mask R-CNN in Cai & Vasconcelos (2017). Hyperparameters used for DARL pre-trained with diffusion loss are listed in Table 11.
Table 11: Hyperparameters for COCO Experiments
Name Value
Resolution 1024 x 1024
Drop path 0.4
Augmentation panoptic deeplab policy †
Random scaling [0.1, 2.5]
Learning rate schedule Cosine
Warm-up steps 256
Training steps 73920
Batch size 64
Optimizer AdamW
Optimizer momentum $\beta_{1}$=0.9, $\beta_{2}$=0.999
Learning rate 7.5e-4 (DARL); 5e-5 (MAE/SUP)
Weight decay 0.1
Layer-wise lr decay 0.85 (DARL); 0.8 (MAE/SUP)
Exponential moving average decay 0.9998
@@ A.4 VTAB
A.5 Two-Stream Architecture
Appendix B Further Results and Discussions
B.1 Linear Evaluation
(a) Training Epochs
(b) Network Depth
Figure 8: ImageNet top-1 accuracy with linear evaluation protocol. 8(a)) Linear performance increases with longer training schedule. Models are ViT-L16 pre-trained with diffusion objective. 8(b)) Linear performance varies with the layer depth. Layer depth is normalized to [0, 1]. Models are pre-trained with diffusion objective.
@@ B.1 Linear Evaluation
B.2 Normalized Pixels
He et al. (2021) introduced normalized pixels as target of the reconstruction. More precisely, instead of using the raw or standardized pixels, they use pixel values normalized by the mean and standard deviation of the patch statistics as the target. He et al. (2021) and Wei et al. (2023) find that using the patch normalized pixel value improves the quality of the representation. We experiment normalized pixels with DARL + MSE loss. However, Table 14 suggests using normalized pixels doesn’t improve the performance in our case.
Table 14: ImageNet Top-1 Accuracy using Normalized Pixels as Target
Architecture
Target ViT-B16 ViT-L16 ViT-H14
Pixels 82.7 84.7 85.5
Norm. Pixels 82.4 84.5 85.5
All the models are pre-trained for 800 epochs and fine-tuned for 90 epochs.
B.3 Scaling Behavior
While we don’t have a systematic study of models with different sizes beyond the standard ViT family, for the standard model sizes we trained, we observe a correlation between validation loss during pre-training and downstream performance for both MSE and diffusion loss of the same noise schedule (see Table 15).
Table 15: Validation Loss v.s. ImageNet Accuracy
MSE DIFF †
Loss Top-1 Acc (%) Loss Top-1 Acc (%)
0.194 83.6 0.179 83.5
0.189 84.2 0.174 84.3
0.185 84.5 0.171 84.8
0.182 84.7 0.168 84.9
0.158 85.5 0.146 85.9
†Diffusion loss with beta distribution noise schedule of a=0.03, b=1.
While there is not enough datapoints to make a statistically significant analysis for the moment, we believe this is a viable direction for future research enabled by our framework.
B.4 Tokenization
One area to further explore is the use of a tokenizer. We provide some preliminary results using a discrete VAE encoder (same tokenizer as in Dall-E and BeiT). The tokenizer serves two functions: encoding the image patch to form the inputs of the Transformer; serving as a target for the outputs of the Transformer.
Table 16: ImageNet Top-1 Accuracy with dVAE Tokenzier
Supervised Unsupervised
Default dVAE inputs MSE Denoising dVAE input+MSE dVAE target
82.7 75.9 83.6 83.5 62.9 82.9
Supervised uses non-causal Transformer. The default model uses linear patch embedding.
Unsupervised uses DARL pre-trained for 100 epochs. The default model uses linear patch embedding and denosing patch decoder.
Table 16 shows the top-1 accuracy on ImageNet. Supervised model is ViT-L16 without causal attention masking and uses 2D RoPE. Unsupervised model is DARL with ViT-L16 backbone pre-trained 100 epochs. Both MSE and denoising objectives perform similarly. We can see that using dVAE to encode the image patches has a significant negative impact on model performance, both in supervised and unsupervised. However, dVAE as the target is only slightly worse than MSE or denoising. It worth further investigating for longer training schedule and evaluate on other downstream tasks.
B.5 Discussion on Noise Schedule
Figure 9: Comparison of noise schedules between DDPM, l-DAE and DARL.
Figure 9 compares our schedule with both the original DDPM schedule and the linear decay schedule proposed by Chen et al. (2024). We could see that DDPM schedule samples 80% from values [0.1, 0.6] and with a slight peak around 0.3. Linear decay schedule samples mostly from very small noise regions. Optimal schedule for DARL samples mainly from high noise regions. As we discussed in the main text, the difference in the preferred noise schedule between l-DAE and our work can be understood as due to the different tasks imposed by the framework. The representation learning task of l-DAE is denoising. Therefore, totally destroying the image is problematic. In our case, we combine the autoregressive prediction with denoising.
@@ Appendix C Visualizations for Patch Ordering
Figure 10 compares the patch reconstruction error at each postion in the sequence between different fixed order strategies. Figure 11 and Figure 12 show the visualization and the average per-patch reconstruction errors for nested raster order. Figure 13 and Figure 14 show the visualization and the average per-patch reconstruction errors for round-robin order. See Section 4.4 for details.
@@ Appendix E Diffusion
We use the mathematical framework established by DDPM (Ho et al., 2020). Many of the analysis can be found in DDPM paper. We present here for the completeness of our analysis. To simplify notation, we use $x$ as the image or individual patch and subscript $t$ for the diffusion process.